OpenStudy (anonymous):

find the value of the investment 20 years from now. the equation is f(t)=4000(2)^t/6) or f(t)=1.1225^t either one you can use.please help me

OpenStudy (anonymous):

can i get yo number though?

OpenStudy (tkhunny):

Have you considered substituting t = 20 into your equations?

OpenStudy (anonymous):

you want my number?

OpenStudy (anonymous):

naw just jk but u cute do

OpenStudy (anonymous):

and yes i know that i have to do that but i cant seem to get the answer from like the book that is where i am lost

OpenStudy (anonymous):

i thought u ment it lol and thanks.

OpenStudy (anonymous):

lol the only way you can find intelligent girls is on openstudy haha

OpenStudy (tkhunny):

Please share the book's answer. Something fishy so far. Also, it would help to have the EXACT problem statement. Did it say something like, "An investment doubles every six years"?

OpenStudy (anonymous):

this is what is says a person invests $4000 in stocks today, and their value doubles every 6 years. let f(t) be the value ( in dollars) of the investment at t years from now.

OpenStudy (anonymous):

why do you say that you can only find intelligent girls on here only?

OpenStudy (tkhunny):

And the book doesn't like $40,317.47?

OpenStudy (anonymous):

yeah that is the answer but i cant seem to get that answer

OpenStudy (tkhunny):

Well then, you'll have to show me what buttons you are pushing on your calculator. I just used your formula.

OpenStudy (anonymous):

i was doing it step by step on paper

OpenStudy (tkhunny):

Hmmm... 1.1225^20 is a bit tedious with a pencil. Anyway, you have the right idea.

OpenStudy (anonymous):

so is it better to do on the calculator? im just lost with how i am suppose to do this. before i was just working on each side.

OpenStudy (tkhunny):

Well, they are faster and more accurate. Let's see. 20/6 = 10/3 = 3 + 1/3 It looks like your need the cube root of two, whatever you do. You'll need a calculator. It doesn't have to be an expensive one. Just make sure it has an exponent button, usually something like \[x^{y}\]

OpenStudy (anonymous):

thanks.